What is risk management strategy in term of how much should a trader trade (i.e. risk per trade) and stop loss.
Answer #1

Every trader must have a risk management strategy in place even before placing his first trade.

Generally, a trader should not risk more than 2% of his capital in any single trade. This is not the amount invested in a trade but this is the amount trader may lose if the trade goes against him, that is when the stop loss is hit.

Let's consider an example. The traders total capital is $100,000. The trader has a bullish view on a stock that is trading at $80 per share and he thinks the stock may go to $84. He keeps a stop loss of $78. That is he is risking $2 to make $4.

Now, how many shares to buy depends on this stop loss. He loses $2 if the stock drops. He is ready to risk 2% of his entire capital. That is 100,000 * 0.02 = $2,000

Number of shares to buy = 2,000/2 = 1,000 shares.

If he buys 1,000 shares and the stock hits stop loss, he is going to lose $2,000, which is 2% of his capital. Of course, sometimes he may lose more than $2 per share due to gap down in stock price. But, generally, this is one way of managing risk.

If the trader is a novice, he can choose to risk only 1% of his capital per trade. In that case, with the same stop loss, he would buy only 500 shares. When the stop loss hits he would lose $1,000, which is 1% of his capital.

Can you please upvote? Thank You :-)